# The extension of Schwarzschild line element to include uniformly accelerated mass

### v3

@@ 3 Metric for a uniformly accelerated Schwarzschild object
$ds^{2}$ $=\left[1-\frac{2m}{\sqrt{\bar{x}^{2}+\bar{y}^{2}+\bar{z}^{2}}}\right]c^{2}d\bar{t}^{2}-\Bigg{[}\Bigg{\{}\left(\bar{x}^{2}+\bar{y}^{2}+\bar{z}^{2}\right)^{3/2}-2m\left(\bar{y}^{2}+\bar{z}^{2}\right)\Bigg{\}}d\bar{x}^{2}$
$-\Bigg{\{}\Big{(}\bar{x}^{2}+\bar{y}^{2}+\bar{z}^{2}\Big{)}^{3/2}-2m\left(\bar{x}^{2}+\bar{z}^{2}\right)\Bigg{\}}\bar{y}^{2}-\Bigg{\{}\left(\bar{x}^{2}+\bar{y}^{2}+\bar{z}^{2}\right)^{3/2}$
$-2m\left(\bar{x}^{2}+\bar{y}^{2}\right)\Bigg{\}}d\bar{z}^{2}-4m\bar{x}\bar{y}d\bar{x}\,d\bar{y}-4m\bar{y}\bar{z}d\bar{y}\,d\bar{z}-4m\bar{x}\bar{z}d\bar{z}\,d\bar{x}\Bigg{]}$
$\times\Big{\{}\left(\bar{x}^{2}+\bar{y}^{2}+\bar{z}^{2}\right)\left(\sqrt{\bar{x}^{2}+\bar{y}^{2}+\bar{z}^{2}}-2m\right)\Big{\}}^{-1}$ (16)
@@
where in order to make the space to be Minkowskian one must have the derivative of the function $V$ with respect to $X$ to be proportional to $U$. The Rindler line element as used in the present work passes this test. But the Schwarzschild line element does not pass this test primarily because it can not be expressed in the form of the above equation (17), due to the presence of cross terms involving $d\bar{x}$ versus $d\bar{y}$ and $d\bar{z}$. However, if we express our Schwarzschild line element in isotropic form under far field approximation such a possibility can be explored.
@@
$ds^{2}$ $=\Bigg{[}\frac{c^{4}\left(1-\frac{2m}{\sqrt{\Delta_{0}}}\right)}{a_{x}^{2}\left(\frac{c^{2}}{a_{x}}+x\right)^{2}\left(\frac{a_{x}^{2}c^{2}t^{2}}{\left(a_{x}x+c^{2}\right)^{2}}-1\right)^{2}}+\frac{c^{2}t^{2}\left(2m\left(y^{2}+z^{2}\right)-\Delta_{0}^{3/2}\right)}{\zeta\Delta_{0}\left(\sqrt{\Delta_{0}}-2m\right)}\Bigg{]}c^{2}dt^{2}$
$+\Bigg{[}\frac{c^{6}t^{2}\left(1-\frac{2m}{\sqrt{\Delta_{0}}}\right)}{a_{x}^{2}\left(\frac{c^{2}}{a_{x}}+x\right)^{4}\left\{1-\frac{a_{x}^{2}c^{2}t^{2}}{\left(a_{x}x+c^{2}\right)^{2}}\right\}^{2}}+\frac{\left(\frac{c^{2}}{a_{x}}+x\right)^{2}\left(2m\left(y^{2}+z^{2}\right)-\Delta_{0}^{3/2}\right)}{\zeta\Delta_{0}\left(\sqrt{\Delta_{0}}-2m\right)}\Bigg{]}dx^{2}$
$+\frac{2m\left(\Delta_{2}\right)-\Delta_{0}^{3/2}}{\Delta_{0}\left(\sqrt{\Delta_{0}}-2m\right)}dy^{2}+\frac{2m\left(\Delta_{1}\right)-\Delta_{0}^{3/2}}{\Delta_{0}\left(\sqrt{\Delta_{0}}-2m\right)}dz^{2}$
$-2\Bigg{[}\frac{c^{5}\,t\,\left(\frac{c^{2}}{a_{x}}+x\right)\left(1-\frac{2m}{\sqrt{\Delta_{0}}}\right)}{a_{x}^{2}\left(\zeta\right)^{2}}+\frac{ct\left(\frac{c^{2}}{a_{x}}+x\right)\left(2m\left(y^{2}+z^{2}\right)-\Delta_{0}^{3/2}\right)}{\zeta\Delta_{0}\left(\sqrt{\Delta_{0}}-2m\right)}\Bigg{]}cdtdx$
$+\frac{4mc^{2}ty\left(\sqrt{\zeta}-\frac{c^{2}}{a_{x}}\right)}{\sqrt{\zeta}\,\Delta_{0}\left(\sqrt{\Delta_{0}}-2m\right)}dtdy+\frac{4mc^{2}tz\left(\sqrt{\zeta}-\frac{c^{2}}{a_{x}}\right)}{\sqrt{\zeta}\,\Delta_{0}\left(\sqrt{\Delta_{0}}-2m\right)}dtdz$
$-\frac{4my\left(\frac{c^{2}}{a_{x}}+x\right)\left(\sqrt{\zeta}-\frac{c^{2}}{a_{x}}\right)}{\sqrt{\zeta}\,\Delta_{0}\left(\sqrt{\Delta_{0}}-2m\right)}dxdy-\frac{4mz\left(\frac{c^{2}}{a_{x}}+x\right)\left(\sqrt{\zeta}-\frac{c^{2}}{a_{x}}\right)}{\sqrt{\zeta}\,\Delta_{0}\left(\sqrt{\Delta_{0}}-2m\right)}dxdz$
$-\frac{4myz}{\Delta_{0}\left(\sqrt{\Delta_{0}}-2m\right)}dydz$ (18)
@@
$\Delta_{0}$ $=\left(\sqrt{\zeta}-\frac{c^{2}}{a_{x}}\right)^{2}+y^{2}+z^{2},\hskip 11.38092pt\Delta_{1}=\left(\sqrt{\zeta}-\frac{c^{2}}{a_{x}}\right)^{2}+y^{2},$ (19)
$\Delta_{2}$ $=\left(\sqrt{\zeta}-\frac{c^{2}}{a_{x}}\right)^{2}+z^{2}\text{\,\,\,and\,\,}\zeta=\left(\frac{c^{2}}{a_{x}}+x\right)^{2}-c^{2}t^{2}$
@@
Further, if we consider that the direction of the proper acceleration $a$ of an accelerated frame is in some arbitrary direction on the $XY$ plane, then it will have two components, one along the $X$ and other, along the $Y$ direction. So, as the acceleration is in some arbitrary direction on the $XY$ plane then the inverse Rindler transformation relations in such a case can be written as (see Appendix B.2) :
@@
$\bar{x}$ $=x+\Bigg{[}-\frac{1}{a_{xy}^{2}}\left(x\,a_{x}+y\,a_{y}\right)+\Bigg{\{}\left(\frac{1}{a_{xy}^{2}}\left(x\,a_{x}+y\,a_{y}\right)+\frac{c^{2}}{a_{xy}^{2}}\right)^{2}-\frac{c^{2}t^{2}}{a_{xy}^{2}}\Bigg{\}}^{1/2}-\frac{c^{2}}{a_{xy}^{2}}\Bigg{]}a_{x}$
$\bar{y}$ $=y+\Bigg{[}-\frac{1}{a_{xy}^{2}}\left(x\,a_{x}+y\,a_{y}\right)+\Bigg{\{}\left(\frac{1}{a_{xy}^{2}}\left(x\,a_{x}+y\,a_{y}\right)+\frac{c^{2}}{a_{xy}^{2}}\right)^{2}-\frac{c^{2}t^{2}}{a_{xy}^{2}}\Bigg{\}}^{1/2}-\frac{c^{2}}{a_{xy}^{2}}\Bigg{]}a_{y}$
@@
The gravitational field due to an accelerated Schwarzschild object moving along XX–direction can be described using the transformation equation (15). However, the transformation relations between inertial (rest frame) and accelerated (non-inertial) frames with non-zero initial velocity can written as follows (see Appendix B.3) [19]:
@@
Furthermore, the inverse transformation relation of (21) can be written as follows (see Appendix B.3):
@@
$ds^{2}$ $=\Bigg{[}\frac{c^{4}\left(a_{x}x+\gamma c^{2}\right)^{2}\left(\sqrt{\xi}-2m\right)}{\Big{\{}a_{x}^{2}\left(c^{2}t^{2}-x^{2}\right)+2a_{x}c\gamma(ctv_{x}-cx)+c^{2}\gamma^{2}\left(v_{x}^{2}-c^{2}\right)\Big{\}}^{2}\sqrt{\xi}}$
$+\frac{\left(\gamma\frac{cv_{x}}{a_{x}}+ct\right)^{2}\left\{2m\left(y^{2}+z^{2}\right)-\xi^{3/2}\right\}}{\left\{\Psi\right\}\xi\left(\sqrt{\xi}-2m\right)}\Bigg{]}c^{2}dt^{2}$
$+\Bigg{[}\frac{c^{4}(a_{x}ct+c\gamma v_{x})^{2}\left(\sqrt{\xi}-2m\right)}{\Big{\{}a_{x}^{2}\left(c^{2}t^{2}-x^{2}\right)+2a_{x}c\gamma(ctv_{x}-cx)+c^{2}\gamma^{2}\left(v_{x}^{2}-c^{2}\right)\Big{\}}^{2}\sqrt{\xi}}$
$+\frac{\left(\gamma\frac{c^{2}}{a_{x}}+x\right)^{2}\left\{2m\left(y^{2}+z^{2}\right)-\xi^{3/2}\right\}}{\left\{\Psi\right\}\xi\left(\sqrt{\xi}-2m\right)}\Bigg{]}dx^{2}$
$+\frac{1}{\xi\left(\sqrt{\xi}-2m\right)}\Bigg{[}2m\left\{\left(\sqrt{\Psi}-\frac{c^{2}}{a_{x}}\right)^{2}+z^{2}\right\}-\xi^{3/2}\Bigg{]}\,\,dy^{2}$
$+\frac{1}{\xi\left(\sqrt{\xi}-2m\right)}\Bigg{[}2m\left\{\left(\sqrt{\Psi}-\frac{c^{2}}{a_{x}}\right)^{2}+y^{2}\right\}-\xi^{3/2}\Bigg{]}\,\,dz^{2}$
$+2\Bigg{[}-\frac{c^{4}\left(a_{x}x+\gamma c^{2}\right)(a_{x}ct+c\gamma v_{x})\left(\sqrt{\xi}-2m\right)}{\Big{\{}a_{x}^{2}\left(c^{2}t^{2}-x^{2}\right)+2a_{x}c\gamma(ctv_{x}-cx)+c^{2}\gamma^{2}\left(v_{x}^{2}-c^{2}\right)\Big{\}}^{2}\sqrt{\xi}}$
$-\frac{\left(\gamma\frac{c^{2}}{a_{x}}+x\right)\left(\gamma\frac{cv_{x}}{a_{x}}+ct\right)\left\{2m\left(y^{2}+z^{2}\right)-\xi^{3/2}\right\}}{\left\{\Psi\right\}\xi\left(\sqrt{\xi}-2m\right)}\Bigg{]}\,\,cdxdt$
$+\Bigg{[}\frac{4my\left(\gamma\frac{cv_{x}}{a_{x}}+ct\right)\left(\sqrt{\Psi}-\frac{c^{2}}{a_{x}}\right)}{\sqrt{\Psi}\xi\left(\sqrt{\xi}-2m\right)}\Bigg{]}\,cdy\,dt+\Bigg{[}\frac{4mz\left(\gamma\frac{cv_{x}}{a_{x}}+ct\right)\left(\sqrt{\Psi}-\frac{c^{2}}{a_{x}}\right)}{\sqrt{\Psi}\xi\left(\sqrt{\xi}-2m\right)}\Bigg{]}\,cdzdt$
$+\Bigg{[}-\frac{4my\left(\gamma\frac{c^{2}}{a_{x}}+x\right)\left(\sqrt{\Psi}-\frac{c^{2}}{a_{x}}\right)}{\sqrt{\Psi}\xi\left(\sqrt{\xi}-2m\right)}\Bigg{]}\,dxdy+\Bigg{[}-\frac{4mz\left(\gamma\frac{c^{2}}{a_{x}}+x\right)\left(\sqrt{\Psi}-\frac{c^{2}}{a_{x}}\right)}{\sqrt{\Psi}\xi\left(\sqrt{\xi}-2m\right)}\Bigg{]}\,dzdx$
$+\Bigg{[}-\frac{4myz}{\xi\left(\sqrt{\xi}-2m\right)}\Bigg{]}\,dy\,dz$ (23)
@@
$\xi$ $=\left(\sqrt{\Psi}-\frac{c^{2}}{a_{x}}\right)^{2}+y^{2}+z^{2}\text{\,\,\,and\,\,\,\,}\Psi=\left(\gamma\frac{c^{2}}{a_{x}}+x\right)^{2}-\left(\gamma\frac{cv_{x}}{a_{x}}+ct\right)^{2}$ (24)
@@
    $ds^{2}$ $=\gamma^{2}\Bigg{[}1-\frac{2m}{\sqrt{\mathcal{R}}}+\frac{v_{x}^{2}}{c^{2}}\Bigg{\{}2m\left(y^{2}+z^{2}\right)-\Big{(}\mathcal{R}\Big{)}^{3/2}\Bigg{\}}\Bigg{\{}\Big{(}\mathcal{R}\Big{)}\Big{(}\sqrt{\mathcal{R}}-2m\Big{)}\Bigg{\}}^{-1}\Bigg{]}c^{2}dt^{2}$
    $+\gamma^{2}\Bigg{[}\frac{2m\left(y^{2}+z^{2}\right)-\left(\mathcal{R}\right)^{3/2}}{\left(\mathcal{R}\right)\left(\sqrt{\mathcal{R}}-2m\right)}+\frac{v_{x}^{2}}{c^{2}}\left(1-\frac{2m}{\sqrt{\mathcal{R}}}\right)\Bigg{]}dx^{2}$
    $+\Bigg{[}\frac{2m\left(\gamma^{2}(x-v_{x}t)^{2}+z^{2}\right)-\left(\mathcal{R}\right)^{3/2}}{\left(\mathcal{R}\right)\left(\sqrt{\mathcal{R}}-2m\right)}\Bigg{]}dy^{2}+\Bigg{[}\frac{2m\left(\gamma^{2}(x-v_{x}t)^{2}+y^{2}\right)-\left(\mathcal{R}\right)^{3/2}}{\left(\mathcal{R}\right)\left(\sqrt{\mathcal{R}}-2m\right)}\Bigg{]}dz^{2}$
    $+\frac{4\gamma^{2}m\,v_{x}\Bigg{\{}\sqrt{\mathcal{R}}\left(\gamma^{2}(x-v_{x}t)^{2}+\mathcal{R}\right)-2m\left(\mathcal{R}\right)\Bigg{\}}}{\left(\mathcal{R}\right)^{3/2}\left(\sqrt{\mathcal{R}}-2m\right)}dx\,dt$
    $+\Bigg{[}\frac{4\gamma^{2}my\,v_{x}(x-v_{x}t)}{\left(\mathcal{R}\right)\left(\sqrt{\mathcal{R}}-2m\right)}\Bigg{]}\,dydt+\Bigg{[}\frac{4\gamma^{2}mz\,v_{x}(x-v_{x}t)}{\left(\mathcal{R}\right)\left(\sqrt{\mathcal{R}}-2m\right)}\Bigg{]}dz\,dt$
    $-\Bigg{[}\frac{4\gamma^{2}my(x-v_{x}t)}{\left(\mathcal{R}\right)\left(\sqrt{\mathcal{R}}-2m\right)}\Bigg{]}\,dydx-\Bigg{[}\frac{4myz}{\left(\mathcal{R}\right)\left(\sqrt{\mathcal{R}}-2m\right)}\Bigg{]}dy\,dz$
    $-\Bigg{[}\frac{4\gamma^{2}mz(x-v_{x}t)}{\left(\mathcal{R}\right)\left(\sqrt{\mathcal{R}}-2m\right)}\Bigg{]}dz\,dx$ (25)
@@
    $ds^{2}$ $=\Bigg{[}\frac{c^{4}\left(1-\frac{2m}{\sqrt{\Delta_{0}}}\right)}{a_{x}^{2}\left(\frac{c^{2}}{a_{x}}+x\right)^{2}\left(\frac{a_{x}^{2}c^{2}t^{2}}{\left(a_{x}x+c^{2}\right)^{2}}-1\right)^{2}}+\frac{c^{2}t^{2}\left(2m\left(y^{2}+z^{2}\right)-\Delta_{0}^{3/2}\right)}{\zeta\Delta_{0}\left(\sqrt{\Delta_{0}}-2m\right)}\Bigg{]}c^{2}dt^{2}$
    $+\Bigg{[}\frac{c^{6}t^{2}\left(1-\frac{2m}{\sqrt{\Delta_{0}}}\right)}{a_{x}^{2}\left(\frac{c^{2}}{a_{x}}+x\right)^{4}\left\{1-\frac{a_{x}^{2}c^{2}t^{2}}{\left(a_{x}x+c^{2}\right)^{2}}\right\}^{2}}+\frac{\left(\frac{c^{2}}{a_{x}}+x\right)^{2}\left(2m\left(y^{2}+z^{2}\right)-\Delta_{0}^{3/2}\right)}{\zeta\Delta_{0}\left(\sqrt{\Delta_{0}}-2m\right)}\Bigg{]}dx^{2}$
    $+\frac{2m\left(\Delta_{2}\right)-\Delta_{0}^{3/2}}{\Delta_{0}\left(\sqrt{\Delta_{0}}-2m\right)}dy^{2}+\frac{2m\left(\Delta_{1}\right)-\Delta_{0}^{3/2}}{\Delta_{0}\left(\sqrt{\Delta_{0}}-2m\right)}dz^{2}$
    $-2\Bigg{[}\left(1-\frac{2m}{\sqrt{\Delta_{0}}}\right)\frac{c^{5}\,t\,\left(\frac{c^{2}}{a_{x}}+x\right)}{a_{x}^{2}\left(\zeta\right)^{2}}+\frac{ct\left(\frac{c^{2}}{a_{x}}+x\right)\left(2m\left(y^{2}+z^{2}\right)-\Delta_{0}^{3/2}\right)}{\zeta\Delta_{0}\left(\sqrt{\Delta_{0}}-2m\right)}\Bigg{]}cdtdx$
    $+\frac{4mcty\left(\sqrt{\zeta}-\frac{c^{2}}{a_{x}}\right)}{\sqrt{\zeta}\Delta_{0}\left(\sqrt{\Delta_{0}}-2m\right)}cdtdy+\frac{4mctz\left(\sqrt{\zeta}-\frac{c^{2}}{a_{x}}\right)}{\sqrt{\zeta}\Delta_{0}\left(\sqrt{\Delta_{0}}-2m\right)}cdtdz$
    $-\frac{4my\left(\frac{c^{2}}{a_{x}}+x\right)\left(\sqrt{\zeta}-\frac{c^{2}}{a_{x}}\right)}{\sqrt{\zeta}\Delta_{0}\left(\sqrt{\Delta_{0}}-2m\right)}dxdy-\frac{4mz\left(\frac{c^{2}}{a_{x}}+x\right)\left(\sqrt{\zeta}-\frac{c^{2}}{a_{x}}\right)}{\sqrt{\zeta}\Delta_{0}\left(\sqrt{\Delta_{0}}-2m\right)}dxdz$
    $-\frac{4myz}{\Delta_{0}\left(\sqrt{\Delta_{0}}-2m\right)}dydz$ (26)
@@
    $ds^{2}$ $=\left[1-\frac{2m}{\sqrt{x^{2}+y^{2}+z^{2}}}\right]c^{2}dt^{2}-\Bigg{[}\Bigg{\{}\left(x^{2}+y^{2}+z^{2}\right)^{3/2}-2m\left(y^{2}+z^{2}\right)\Bigg{\}}dx^{2}$
    $-\Bigg{\{}\Big{(}x^{2}+y^{2}+z^{2}\Big{)}^{3/2}-2m\left(x^{2}+z^{2}\right)\Bigg{\}}dy^{2}-\Bigg{\{}\left(x^{2}+y^{2}+z^{2}\right)^{3/2}$
    $-2m\left(x^{2}+y^{2}\right)\Bigg{\}}dz^{2}-4mxydx\,dy-4myzdy\,dz-4mxzdz\,dx\Bigg{]}$
    $\times\Big{\{}\left(x^{2}+y^{2}+z^{2}\right)\left(\sqrt{x^{2}+y^{2}+z^{2}}-2m\right)\Big{\}}^{-1}$ (27)
    which is exactly the Schwarzschild metric in Cartesian coordinates shown in (4) (since $ct=ct^{\prime}$, $x=x^{\prime}$, $y=y^{\prime}$ and $z=z^{\prime}$ for $v_{x}\longrightarrow 0$ and $a_{x}\longrightarrow 0$).
@@
    $ds^{2}$ $=\Bigg{[}\frac{c^{4}\left(x+\frac{c^{2}}{a_{x}}\right)^{2}}{a_{x}^{2}\,\zeta^{2}}-\frac{c^{2}t^{2}}{\zeta}\Bigg{]}(cdt)^{2}-\Bigg{[}\frac{c^{6}t^{2}}{a_{x}^{2}\,\zeta^{2}}-\frac{\left(\frac{c^{2}}{a_{x}}+x\right)^{2}}{\zeta}\Bigg{]}(dx)^{2}-(dy)^{2}$
    $-(dz)^{2}+2\Bigg{[}-\frac{ct}{a_{x}^{2}\,\zeta^{2}}\left(a_{x}x+c^{2}\right)\Big{\{}a_{x}\left(c^{2}t^{2}-x^{2}\right)-2c^{2}x\Big{\}}\Bigg{]}cdtdx$ (28)
@@
The transformation relation for an accelerated frame with respect to an inertial frame $K$ where the acceleration is in arbitrary direction on the $XY$ plane having zero initial velocity is shown in equation (20) (also see Appendix B.2). So using this transformation relation in equation (16) we obtain the line element for a Schwarzschild object having acceleration in some arbitrary direction on the $XY$ plane as below:
$ds^{2}$ $=g_{t\,t}c^{2}dt^{2}+g_{x\,x}dx^{2}+g_{y\,y}dy^{2}+g_{z\,z}dz^{2}+2\Big{\{}g_{x\,t}dxdt+g_{y\,t}dydt$
$+g_{z\,t}dzdt+g_{x\,y}dxdy+g_{y\,z}dydz+g_{z\,x}dzdx\Big{\}}$ (29)

### v4

@@ 4 Metric for a uniformly accelerated Schwarzschild object
$ds^{2}=\left(1-\frac{2m}{\bar{r}-m}\right)c^{2}d\bar{t}^{2}-\left(1+\frac{2m}{\bar{r}-m}\right)\Big\{d\bar{x}^{2}+d\bar{y}^{2}+d\bar{z}^{2}\Big\}$ (21)
where $\bar{r}=\sqrt{\bar{x}^{2}+\bar{y}^{2}+\bar{z}^{2}}$.
@@
where in order to make the space to be Minkowskian one must have the derivative of the function $V$ with respect to $X$ to be proportional to $U$. The Rindler line element as used in the present work passes this test. Thus, as stated before in the case of uniform motion, the Rindler transformation also cannot be used directly to the standard Schwarzschild line element. However, in our present case we assume a far-field approximation such that the spacetime geometry is close to flat, and hence we can use the said transformation relations on the approximated line element (21).
@@
$ds^{2}$ $=\Bigg[\frac{c^{4}}{a_{x}^{2}\left\{\chi\right\}^{2}}\Big\{1-\frac{2m}{\Delta_{0}}\Big\}\left(\frac{c^{2}}{a_{x}}+x\right)^{2}-\frac{c^{2}t^{2}}{\chi}\left(1+\frac{2m}{\Delta_{0}}\right)\Bigg]c^{2}dt^{2}$
$-\Bigg[\frac{1}{\chi}\left(\frac{c^{2}}{a_{x}}+x\right)^{2}\left(1+\frac{2m}{\Delta_{0}}\right)-\frac{c^{6}t^{2}}{a_{x}^{2}\left\{\chi\right\}^{2}}\left(1-\frac{2m}{\Delta_{0}}\right)\Bigg]dx^{2}$
$-\Bigg[1+\frac{2m}{\Delta_{0}}\Bigg]\Big\{dy^{2}+dz^{2}\Big\}+2\Bigg[\frac{ct}{\chi}\left(\frac{c^{2}}{a_{x}}+x\right)\left(1+\frac{2m}{\Delta_{0}}\right)$
$-\frac{c^{6}t}{a_{x}^{2}\left\{\chi\right\}^{2}}\left(1-\frac{2m}{\Delta_{0}}\right)\left(\frac{c^{2}}{a_{x}}+x\right)\Bigg]cdtdx$ (23)
@@
$\Delta_{0}$ $=\Bigg[\Bigg\{\chi^{1/2}-\frac{c^{2}}{a_{x}}\Bigg\}^{2}+y^{2}+z^{2}\Bigg]^{1/2}-m\,\,\,\,\text{and}\hskip 14.22636pt\chi=\left(\frac{c^{2}}{a_{x}}+x\right)^{2}-c^{2}t^{2}$ (24)
@@
Further, if we consider that the direction of the proper acceleration $a$ of an accelerated frame is in some arbitrary direction on the $XY$ plane, then it will have two components, one along the $X$ and other, along the $Y$ direction. So, as the acceleration is in some arbitrary direction on the $XY$ plane then the inverse Rindler transformation relations in such a case can be written as (see Appendix A.2) :
@@
$\bar{x}$ $=x+\Bigg[-\frac{1}{a_{xy}^{2}}\left(xa_{x}+ya_{y}\right)+\Bigg\{\left(\frac{1}{a_{xy}^{2}}\left(xa_{x}+ya_{y}\right)+\frac{c^{2}}{a_{xy}^{2}}\right)^{2}$
$\hskip 199.16928pt-\frac{c^{2}t^{2}}{a_{xy}^{2}}\Bigg\}^{1/2}-\frac{c^{2}}{a_{xy}^{2}}\Bigg]a_{x}$
$\bar{y}$ $=y+\Bigg[-\frac{1}{a_{xy}^{2}}\left(xa_{x}+ya_{y}\right)+\Bigg\{\left(\frac{1}{a_{xy}^{2}}\left(xa_{x}+ya_{y}\right)+\frac{c^{2}}{a_{xy}^{2}}\right)^{2}$
$\hskip 199.16928pt-\frac{c^{2}t^{2}}{a_{xy}^{2}}\Bigg\}^{1/2}-\frac{c^{2}}{a_{xy}^{2}}\Bigg]a_{y}$
@@
The gravitational field due to an accelerated Schwarzschild object moving along XX–direction can be described using the transformation equation (20). However, the transformation relations between inertial (rest frame) and accelerated (non-inertial) frames with non-zero initial velocity can written as follows (see Appendix A.3) [25]:
@@
Furthermore, the inverse transformation relation of (26) can be written as follows (see Appendix A.3):
@@
$ds^{2}$ $=\Bigg[\frac{c^{4}\left(a_{x}x+\gamma c^{2}\right)^{2}\left(\Delta_{1}-3m\right)}{\left(\Delta_{2}\right)^{2}\left(\Delta_{1}-m\right)}-\frac{\left(\frac{c\gamma v_{x}}{a_{x}}+ct\right)^{2}\left(1+\frac{2m}{\Delta_{1}-m}\right)}{\left(\frac{\gamma c^{2}}{a_{x}}+x\right)^{2}-\left(\frac{c\gamma v_{x}}{a_{x}}+ct\right)^{2}}\Bigg]c^{2}dt^{2}$
$-\Bigg[\frac{\left(\frac{\gamma c^{2}}{a}+x\right)^{2}\left(1+\frac{2m}{\Delta_{1}-m}\right)}{\left(\frac{\gamma c^{2}}{a_{x}}+x\right)^{2}-\left(\frac{c\gamma v_{x}}{a_{x}}+ct\right)^{2}}-\frac{c^{4}(a_{x}ct+c\gamma v_{x})^{2}\left(\Delta_{1}-3m\right)}{\left(\Delta_{2}\right)^{2}\left(\Delta_{1}-m\right)}\Bigg]dx^{2}$
$-\Bigg[1+\frac{2m}{\Delta_{1}-m}\Bigg]\Big\{dy^{2}+dz^{2}\Big\}+2\Bigg[\frac{\left(\frac{\gamma c^{2}}{a_{x}}+x\right)\left(\frac{c\gamma v_{x}}{a_{x}}+ct\right)\left(1+\frac{2m}{\Delta_{1}-m}\right)}{\left(\frac{\gamma c^{2}}{a_{x}}+x\right)^{2}-\left(\frac{c\gamma v_{x}}{a_{x}}+ct\right)^{2}}$
$-\frac{c^{4}\left(a_{x}x+\gamma c^{2}\right)(a_{x}ct+c\gamma v_{x})\left(\Delta_{1}-3m\right)}{\left(\Delta_{2}\right)^{2}\left(\Delta_{1}-m\right)}\Bigg]cdtdx$ (28)
@@
$\Delta_{1}$ $=\Bigg\{\Bigg\{\Bigg\{\left(\frac{\gamma c^{2}}{a_{x}}+x\right)^{2}-\left(\frac{c\gamma v_{x}}{a_{x}}+ct\right)^{2}\Bigg\}^{1/2}-\frac{c^{2}}{a_{x}}\Bigg\}^{2}+y^{2}+z^{2}\Bigg\}^{1/2}$ (29)
$\Delta_{2}$ $=a_{x}^{2}\left(c^{2}t^{2}-x^{2}\right)+2a_{x}c\gamma(ctv_{x}-cx)+c^{2}\gamma^{2}\left(v_{x}^{2}-c^{2}\right)$ (30)
@@
    $ds^{2}$ $=\Bigg[1-\frac{2m\gamma^{2}}{\mathcal{R}}\Big\{1+\frac{v_{x}^{2}}{c^{2}}\Big\}\Bigg]c^{2}dt^{2}-\Bigg[1+\frac{2m\gamma^{2}}{\mathcal{R}}\Big\{1+\frac{v_{x}^{2}}{c^{2}}\Big\}\Bigg]dx^{2}$
    $-\Big[1+\frac{2m}{\mathcal{R}}\Big]\Big\{dy^{2}+dz^{2}\Big\}+\frac{8\gamma^{2}mv_{x}}{\mathcal{R}}\,dxdt$ (31)
@@
    $ds^{2}$ $=\Bigg[\frac{c^{4}}{a_{x}^{2}\left\{\chi\right\}^{2}}\Big\{1-\frac{2m}{\Delta_{0}}\Big\}\left(\frac{c^{2}}{a_{x}}+x\right)^{2}-\frac{c^{2}t^{2}}{\chi}\left(1+\frac{2m}{\Delta_{0}}\right)\Bigg]c^{2}dt^{2}$
    $-\Bigg[\frac{1}{\chi}\left(\frac{c^{2}}{a_{x}}+x\right)^{2}\left(1+\frac{2m}{\Delta_{0}}\right)-\frac{c^{6}t^{2}}{a_{x}^{2}\left\{\chi\right\}^{2}}\left(1-\frac{2m}{\Delta_{0}}\right)\Bigg]dx^{2}$
    $-\Bigg[1+\frac{2m}{\Delta_{0}}\Bigg]\Big\{dy^{2}+dz^{2}\Big\}+2\Bigg[\frac{ct}{\chi}\left(\frac{c^{2}}{a_{x}}+x\right)\left(1+\frac{2m}{\Delta_{0}}\right)$
    $-\frac{c^{6}t}{a_{x}^{2}\left\{\chi\right\}^{2}}\left(1-\frac{2m}{\Delta_{0}}\right)\left(\frac{c^{2}}{a_{x}}+x\right)\Bigg]cdtdx$ (32)
@@
    $ds^{2}$ $=\left(1-\frac{2m}{r-m}\right)c^{2}dt^{2}-\left(1+\frac{2m}{r-m}\right)\Big\{dx^{2}+dy^{2}+dz^{2}\Big\}$ (33)
    which is exactly the Schwarzschild metric in Cartesian coordinates shown in (8) (since $ct=ct^{\prime}$, $x=x^{\prime}$, $y=y^{\prime}$, $z=z^{\prime}$ and $r=r^{\prime}$ for $v_{x}\longrightarrow 0$ and $a_{x}\longrightarrow 0$).
@@
    $ds^{2}$ $=\Bigg[\frac{c^{4}}{a_{x}^{2}\left\{\chi\right\}^{2}}\left(\frac{c^{2}}{a_{x}}+x\right)^{2}-\frac{c^{2}t^{2}}{\chi}\Bigg]c^{2}dt^{2}-\Bigg[\frac{1}{\chi}\left(\frac{c^{2}}{a_{x}}+x\right)^{2}-\frac{c^{6}t^{2}}{a_{x}^{2}\left\{\chi\right\}^{2}}\Bigg]dx^{2}$
    $-\Big\{dy^{2}+dz^{2}\Big\}+2\Bigg[\frac{ct}{\chi}-\frac{c^{6}t}{a_{x}^{2}\left\{\chi\right\}^{2}}\Bigg]\left(\frac{c^{2}}{a_{x}}+x\right)cdtdx$ (34)
@@
The transformation relation for an accelerated frame with respect to an inertial frame $K$ where the acceleration is in arbitrary direction on the $XY$ plane having zero initial velocity is shown in equation (25) (also see Appendix A.2). So using this transformation relation in equation (21) we obtain the line element for a Schwarzschild object having acceleration in some arbitrary direction on the $XY$ plane as below:
$ds^{2}$ $=g_{t\,t}c^{2}dt^{2}+g_{x\,x}dx^{2}+g_{y\,y}dy^{2}+g_{z\,z}dz^{2}$
$\hskip 28.45274pt+2\Big\{g_{x\,t}dxcdt+g_{y\,t}dycdt+g_{x\,y}dxdy\Big\}$ (35)